OpenStudy (anonymous):

Write the equation of the line perpendicular to the line x - 3y = 9 and passing through the point (3,5).

OpenStudy (anonymous):

@danish071996 @perl

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

do you know the slope of \(x-3y=9\)?

OpenStudy (anonymous):

heck no

OpenStudy (anonymous):

then you can't do it

OpenStudy (anonymous):

i know thats why i got it on here so i can get help haha

OpenStudy (anonymous):

solve for \(y\) to get the slope

OpenStudy (anonymous):

or else memorize that the slope of \[ax+by=c\] is \[-\frac{a}{b}\]

OpenStudy (anonymous):

forget this crap lol

OpenStudy (anonymous):

yeah it is too hard

OpenStudy (anonymous):

so the answer is -3y=-x+9

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

answer is y=3x-4

OpenStudy (anonymous):

wrong